OpenStudy (anonymous):

What is the distance between the points (-3, -10) and (4, 9)?

OpenStudy (anonymous):

Find the difference in the x value and y value...Or use a graph

OpenStudy (alexandervonhumboldt2):

use distance formula: \[\sqrt[2]{(x_1-x_2)^2+(y_1-y_2)^2}\]

OpenStudy (mrnood):

Could I suggest you first plot the 2 points on a chart? That way you can understand what you are measuring. Add the points to this charts: 

OpenStudy (alexandervonhumboldt2):

plug in values and solve: \[\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}=\sqrt{(-3-4)^2+(-10-9)^2}=?\]

OpenStudy (anonymous):

Is the Answer -410?

OpenStudy (mrnood):

a distance cannot be negative in this case

OpenStudy (anonymous):

So a positive 410

OpenStudy (alexandervonhumboldt2):

how you got 410???????????

OpenStudy (mathmale):

Alex poses a good question. Please, show all of your work and explain your reasoning, instead of asking, "is the answer -410?"

OpenStudy (mrnood):

 if you did the sketch like I suggested you can see that 410 is not a sensible answer the points are nowhere near that distance apart